A ball is dropped from a height of 64 feet. Its height above the earth in feet is given by h(t ) = -16t2 + 64, where t is the number of seconds after it is dropped.
a) Find h(1).
b) How long does it take the ball to fall to the earth?

Answer by nerdybill(7384)

A ball is dropped from a height of 64 feet. Its height above the earth in feet is given by h(t ) = -16t^2 + 64, where t is the number of seconds after it is dropped.
a) Find h(1).
Simply substitute t with a 1 and solve:
h(1) = -16(1)^2 + 64
h(1) = -16(1) + 64
h(1) = -16 + 64
h(1) = 48 feet
.
b) How long does it take the ball to fall to the earth?
Simply substitute h(t) with a 0 and solve for t-- height when it falls to earth.
0 = -16t^2 + 64
16t^2 = 64
t^2 = 64/16
t^2 = 4
t = 2 seconds